# Test-Time Adaptation of Vision-Language Models for Open-Vocabulary Semantic Segmentation

### v1

@@ Abstract
Recently, test-time adaptation has attracted wide interest in the context of vision-language models for image classification. However, to the best of our knowledge, the problem is completely overlooked in dense prediction tasks such as Open-Vocabulary Semantic Segmentation (OVSS). In response, we propose a novel TTA method tailored to adapting VLMs for segmentation during test time. Unlike TTA methods for image classification, our Multi-Level and Multi-Prompt (MLMP) entropy minimization integrates features from intermediate vision-encoder layers and is performed with different text-prompt templates at both the global CLS token and local pixel-wise levels. Our approach could be used as plug-and-play for any segmentation network, does not require additional training data or labels, and remains effective even with a single test sample. Furthermore, we introduce a comprehensive OVSS TTA benchmark suite, which integrates a rigorous evaluation protocol, seven segmentation datasets, and 15 common corruptions, with a total of 82 distinct test scenarios, establishing a standardized and comprehensive testbed for future TTA research in open-vocabulary segmentation. Our experiments on this suite demonstrate that our segmentation-tailored method consistently delivers significant gains over direct adoption of TTA classification baselines. Our code can be found at our repository.
@@ 1 Introduction
Although existing OVSS methods have made significant progress, they remain vulnerable to domain shifts at test time, such as environmental changes or image corruptions, which may dramatically degrade segmentation quality. In the absence of a mechanism enabling to adapt to unseen test-time distributions, these models might lose their generalization capabilities, which limits their reliability in real-world applications. Consequently, there is an unresolved gap for the Test-Time Adaptation (TTA) of OVSS models, which would enable models to dynamically adjust both to the task shift of VLM-based segmentation and to the domain shifts encountered during inference.
@@
The core requirement for test-time adaptation is a reliable signal that faithfully reflects the current input distribution—even under severe domain shifts or corruptions. To meet this need, MLMP begins by adaptively integrating intermediate layers of the vision encoder: earlier layers preserve fine-grained edges and textures, while deeper blocks encode semantic context, and each layer reacts differently when the data distribution changes. By aggregating these multi-level features into the adaptation process and weighting them by their confidence, MLMP harvests the most trustworthy signals for each input sample.
Beyond multi-level fusion, MLMP leverages VLMs’ prompt sensitivity to model uncertainty. Prior work watt ; tpt shows that changing a prompt template, e.g., from ‘’a photo of a {class}” to ‘’an origami of a {class}”, could drastically change the classification performance. In segmentation, we show that this effect is even more extreme: per-pixel predictions under different prompts diverge far more than the single CLS token used for classification (Fig. 1a: CLS token vs. last-layer spatial tokens). This sensitivity effect is even more pronounced in intermediate feature maps. The intermediate layers that we fuse for reliability exhibit even stronger template-specific shifts (Fig. 1a: intermediate-layer tokens). Instead of viewing this inconsistency as a weakness, MLMP models it directly by incorporating multi-prompt, multi-level predictions into its adaptation objective function. This multi-prompt approach not only smooths out template-specific noise, thereby reducing gradient variance and preventing degenerate collapse, but also ensures that the model yields segmentations that are consistent across diverse linguistic formulations. In this way, MLMP transforms prompt sensitivity into a powerful adaptation signal that complements its multi-level feature integration. As illustrated in Fig. 1b, our MLMP method consistently outperforms the non-adapted baseline, achieving about 8–9 absolute mIoU improvements on domain-shifted inputs, while also boosting performance on original (non-corrupted) images. This demonstrates the benefit of our joint multi-level, multi-prompt adaptation.
Figure 1: Motivation. (a) Left: Mean $\pm$ std entropy across seven text templates for the CLS token and the spatial tokens of the final and intermediate vision layers. Even the final-layer spatial tokens exhibit higher entropy and variability than CLS, and this sensitivity grows further in intermediate layers (numbers show % std increase relative to CLS). These patterns highlight pronounced prompt-induced uncertainty at multiple depths and motivate both multi-level and multi-prompt adaptation. (b) Right: mIoU of the baseline vs MLMP on clean and corrupted data, showing consistent absolute improvements and underscoring the effectiveness of our joint adaptation strategies.
@@
  • Multi-Prompt Local-Global Test Time Optimization: MLMP turns prompt sensitivity into signal by directly minimizing entropy across different text prompt templates at both the global CLS token and local pixel-wise levels. This optimization naturally complements our mutli-layer feature fusion by enforcing consistency across linguistic perspectives and feature depths.
  • Comprehensive OVSS TTA Benchmark Suite: We curate a rigorous evaluation protocol spanning seven mainstream segmentation datasets and 15 common corruptions, with a total of 82 distinct test scenarios, establishing a standardized and comprehensive testbed for future TTA research in open-vocabulary segmentation. Our experiments on this suite demonstrate that MLMP consistently delivers significant gains over baselines across all scenarios.
@@ 2 Related Work
Test-time adaptation (TTA) for open-vocabulary semantic segmentation (OVSS) remains unexplored—existing TTA methods focus on classification or single-modal segmentation, while OVSS approaches use VLMs without any online adaptation. We bridge this gap with our proposed method MLMP, a plug-and-play TTA framework that could be applied to any OVSS method.
@@
3 Methodology
@@ 3.1 OVSS with VLMs
Following recent approaches for OVSS naclip ; wang2025sclip , we employ a transformer-based VLM to extract visual and text features from the image and concepts in natural language. Specifically, we feed the image $\mathbf{X}$ into the ViT-based vision encoder to extract a visual token matrix $\mathbf{F}=\big{[}\mathbf{f}_{[\texttt{cls}]},\mathbf{f}_{1},\dots,\mathbf{f}_{N}\big{]}$ with each $\mathbf{f}_{i}\in{}^{D}$, where $N\!=\!\lfloor H/s\rfloor\!\times\!\lfloor W/s\rfloor$ is the number of patches of size $s\!\times\!s$ in the image and [cls] is the CLS token for classification. We define $\mathbf{Q}=\big{[}\mathbf{q}_{[\texttt{cls}]},\mathbf{q}_{1},\dots,\mathbf{q}_{N}\big{]}$, with each $\mathbf{q}_{i}\in{}^{D^{\prime}}$, the output features before the projection layer: $\mathbf{F}=\text{proj}(\mathbf{Q})$. At the same time, the text encoder is employed to extract text features $\mathbf{t}_{k}\in{}^{D}$ for each concept $C_{k}\in\mathcal{C}$. This is achieved by combining $C_{k}$ with a text prompt template, for instance “A photo of a [$C_{k}$]” or “An image of a [$C_{k}$]” where $C_{k}$ is an arbitrary text description like “white horse”.
@@
[…]
@@ 3.2 MLMP: Proposed Method
[…]
@@
Uncertainty-Aware Multi-Level Fusion.  In VLM-based classification, the CLS token in the last layer of the visual encoder is typically used to compute the class label probabilities. This approach relies on the idea that the relevant information for classification lies at the end of the ViT and that intermediate layers serve to transform features. In segmentation, however, features from intermediate layers are often used to capture complementary information at different scales unet . This hypothesis is validated in Fig. 1b, and more specifically in Table1), showing that a higher segmentation mIoU is obtained when combining the features from different layers near the end of the ViT.
@@
Image-Level Entropy Minimization.  Since the CLS token is not directly linked to individual patch predictions but rather captures a more global representation of the input, we also include an image-level entropy (ILE) minimization term specifically for this token. As illustrated in Fig. 1, the CLS token demonstrates increased robustness and reliability. This term, which encourages the model to produce more confident global predictions is expressed as:
@@
Multi-Prompt Adaptation.  Prior work on TTA for classification watt has shown the usefulness of leveraging multiple prompt templates in VLM to encode class labels, based on the idea that the templates capture complementary information about these classes. As shown in Fig. 1a, the sensitivity to the choice of prompt templates is even more pronounced in segmentation tasks, where fine-grained spatial predictions are required. Using multiple templates acts as cross-modal regularization, encouraging more stable and generalized learning signals. While different from image augmentation, it can be seen as a strong, safe, and lightweight text-space augmentation. Rather than averaging the weights adapted from different prompt templates as in watt —a computationally expensive approach for dense prediction tasks such as segmentation—our method minimizes our proposed UAML and ILE losses across these templates. Let $\mathbf{T}_{t}$ be the text features obtained using the $t$-th template. Our final adaptation loss is defined as:
@@
A proof of Prop. 1 is provided in Appendix A. ∎
@@ 4 Experimental Settings
Datasets.  In traditional TTA for segmentation, two datasets are commonly employed to simulate domain shifts—one for model training and another for adaptation during inference (e.g., GTAV and Cityscapes)—as both must share the same semantic label space. In our study, as this is the first exploration of TTA for VLMs in segmentation tasks and given that VLMs are pre-trained, we draw inspiration from ImageNet-C cifar10c&10.1 to introduce 15 synthetic corruptions on segmentation datasets. Our experiments are conducted on Pascal VOC 20 (v20), Pascal VOC 21 (v21) PascalVoc , Pascal Context 59 (P59), Pascal Context 60 (P60) PascalContext , and Cityscapes Cityscapes , incorporating both original version (clean) and the synthetic 15 corruptions (denoted with a “-C” suffix). For COCO-Stuff CocoStuff and COCO-Object CocoObject , we use only the original versions. This setup results in 82 distinct test scenarios covering both synthetic and realistic domain shifts, with variations in resolution, scene diversity, object sizes, and class granularity—enabling a comprehensive evaluation of our method’s robustness.
@@ 5.1 Ablation studies
Visualization of Layer-Wise Confidence Weights.  We visualize the mean and standard deviation of the learned layer weights to better understand the behavior of our uncertainty-aware fusion strategy. As shown in Fig. 3, deeper layers tend to receive higher confidence, though earlier layers also contribute, especially under corrupted conditions. This variation is most pronounced in the Cityscapes dataset, where the distribution fluctuates more across layers and corruption types. In contrast, the COCO-Stuff dataset shows a flatter distribution, where the final layer is not consistently the most influential. These results underscore the core strength of our fusion mechanism: its ability to adaptively reweight layers based on input conditions, assigning greater importance to those that remain more reliable under distribution shifts and corruption. Please refer to Appendix H for additional results on other datasets.
@@
V20-C Average 71.92 68.44 72.58
Effect of Number of Prompt Templates.  To isolate this effect, we evaluate performance using different numbers of prompt templates while disabling both the multi-level fusion and the image-level entropy (ILE) term in Eq.10. As shown in Fig.5, increasing the number of templates improves performance up to 7, after which the gains begin to saturate or slightly decline. This trend holds across both clean and corrupted settings, indicating that a moderate number of diverse prompt templates is sufficient for MLMP. We therefore use 7 templates by default in our main experiments. Please refer to Appendix C for template details.
@@
Full Component Analysis.  Table 2 presents an extensive ablation evaluating the contribution of each component in our MLMP strategy. Each proposed element yields consistent gains when added independently, but it is their combination that delivers the highest overall performance across both clean and corrupted settings, highlighting their strong complementary effects within a unified framework. Additionally, we provide further ablations in the Appendix, including alternative OVSS backbones, a different VLM, ViT architecture variants, computational complexity analysis, and MLMP segmentation map visualizations.
Table 3: mIoU comparison of MLMP and baselines across several datasets. CLIPArTT could not be run for a few cases owing to GPU memory shortages. Full per-dataset results are in Appendix J.

### v2

@@ Abstract
Recently, test-time adaptation has attracted wide interest in the context of vision-language models for image classification. However, to the best of our knowledge, the problem is completely overlooked in dense prediction tasks such as Open-Vocabulary Semantic Segmentation (OVSS). In response, we propose a novel TTA method tailored to adapting VLMs for segmentation during test time. Unlike TTA methods for image classification, our Multi-Level and Multi-Prompt (MLMP) entropy minimization integrates features from intermediate vision-encoder layers and is performed with different text-prompt templates at both the global CLS token and local pixel-wise levels. Our approach could be used as plug-and-play for any segmentation network, does not require additional training data or labels, and remains effective even with a single test sample. Furthermore, we introduce a comprehensive OVSS TTA benchmark suite, which integrates a rigorous evaluation protocol, nine segmentation datasets, 15 common synthetic corruptions, and additional real and rendered domain shifts, with a total of 87 distinct test scenarios, establishing a standardized and comprehensive testbed for future TTA research in open-vocabulary segmentation. Our experiments on this suite demonstrate that our segmentation-tailored method consistently delivers significant gains over direct adoption of TTA classification baselines. Code and data are available at https://github.com/dosowiechi/MLMP.
@@ 1 Introduction
Although existing OVSS methods have made significant progress, they remain vulnerable to domain shifts at test time, such as environmental changes or image corruptions, which may dramatically degrade segmentation quality. In the absence of a mechanism enabling them to adapt to unseen test-time distributions, these models might lose their generalization capabilities, which limits their reliability in real-world applications. Consequently, there is an unresolved gap for the Test-Time Adaptation (TTA) of OVSS models, which would enable models to dynamically adjust both to the task shift of VLM-based segmentation and to the domain shifts encountered during inference.
@@
The core requirement for test‐time adaptation is a reliable signal that faithfully reflects the current input distribution—even under severe domain shifts or corruptions. To meet this need, MLMP begins by adaptively integrating intermediate layers of the vision encoder: earlier layers preserve fine‐grained edges and textures, while deeper blocks encode semantic context, and each layer reacts differently when the data distribution changes. By aggregating these multi-level features into the adaptation process and weighting them by their confidence, MLMP harvests the most trustworthy signals for each input sample.
Beyond multi-level fusion, MLMP leverages VLMs’ prompt sensitivity to model uncertainty. Prior work watt; tpt shows that changing a prompt template, e.g., from ‘’a photo of a {class}” to ‘’an origami of a {class}”, could drastically change the classification performance. In segmentation, we show that this effect is even more extreme: per-pixel predictions under different prompts diverge far more than the single CLS token used for classification (Figure 1a: CLS token vs. last-layer spatial tokens). This sensitivity effect is even more pronounced in intermediate feature maps. The intermediate layers that we fuse for reliability exhibit even stronger template-specific shifts (Figure 1a: intermediate-layer tokens). Instead of viewing this inconsistency as a weakness, MLMP models it directly by incorporating multi-prompt, multi-level predictions into its adaptation objective function. This multi-prompt approach not only smooths out template-specific noise, thereby reducing gradient variance and preventing degenerate collapse, but also ensures that the model yields segmentations that are consistent across diverse linguistic formulations. In this way, MLMP transforms prompt sensitivity into a powerful adaptation signal that complements its multi-level feature integration. As illustrated in Figure 1b, our MLMP method consistently outperforms the non-adapted baseline, achieving about 8–9 absolute mIoU improvements on domain-shifted inputs, while also boosting performance on original (non-corrupted) images. This demonstrates the benefit of our joint multi-level, multi-prompt adaptation.
Figure 1: Motivation. (a) Left: Mean $\pm$ std entropy across seven text templates for the CLS token and the spatial tokens of the final and intermediate vision layers. Even the final-layer spatial tokens exhibit higher entropy and variability than CLS, and this sensitivity grows further in intermediate layers (numbers show % std increase relative to CLS). These patterns highlight pronounced prompt-induced uncertainty at multiple depths and motivate both multi-level and multi-prompt adaptation. (b) Right: mIoU of the baseline vs. MLMP on clean and corrupted data, showing consistent absolute improvements and underscoring the effectiveness of our joint adaptation strategies. Here, V20 denotes the Pascal VOC 20 dataset, and V20-C represents the average performance over its 15 synthetic corruption types. The variance in (a) is computed across all samples and all corruptions.
@@
  • Multi-Prompt Local-Global Test-Time Optimization: MLMP turns prompt sensitivity into signal by directly minimizing entropy across different text prompt templates at both the global CLS token and local pixel-wise levels. This optimization naturally complements our multi-level feature fusion by enforcing consistency across linguistic perspectives and feature depths.
  • Comprehensive OVSS TTA Benchmark Suite: We curate a rigorous evaluation protocol spanning nine mainstream segmentation datasets and 15 common synthetic corruptions, and additional real and rendered domain shifts, with a total of 87 distinct test scenarios, establishing a standardized and comprehensive testbed for future TTA research in open-vocabulary segmentation. Our experiments on this suite demonstrate that MLMP consistently delivers significant gains over baselines across all scenarios.
@@ 2 Related Work
Test-time adaptation (TTA) for open-vocabulary semantic segmentation (OVSS) remains unexplored—existing TTA methods focus on classification or single-modality segmentation, while OVSS approaches use VLMs without any online adaptation. We bridge this gap with our proposed method MLMP, a plug-and-play TTA framework that can be applied to any OVSS method.
@@
To better situate this contribution within the broader landscape of general adaptation methods, Table 1 summarizes the key distinctions among zero-shot inference, domain generalization (DG) tfsvit; fds, few-shot segmentation (FSS) panet, test-time training (TTT) tttflow; clust3; NCTTT, and our fully unsupervised test-time adaptation (TTA). This comparison highlights that MLMP addresses the most challenging and realistic scenario, adapting models entirely from unlabeled test data without any source access or annotated support samples.
Table 1: Comparison of general learning and adaptation paradigms. Here, $x^{s},y^{s}$ denote labeled source samples and $x^{t},y^{t}$ denote target (test) samples and labels. Domain Generalization trains on labeled multi-domain source data to improve robustness on unseen targets, while few-shot and test-time training methods rely on labeled or source data during or after training. Our approach (Fully TTA) adapts solely using unlabeled test samples, requiring neither supervision nor source access.
Setting Source Data Target Data Train Loss Test Loss
Zero-Shot Inference $x^{t}$
Domain Generalization (DG) $x^{s},y^{s}$ (often multi-domain) $x^{t}$ $\mathcal{L}(x^{s},y^{s})$ (DG objectives)
Few-Shot Learning (FSS) $x^{t},y^{t}$ (few) $\mathcal{L}(x^{t},y^{t})$ (FSS objectives)
Test-Time Training (TTT) $x^{s},y^{s}$ $x^{t}$ $\mathcal{L}(x^{s},y^{s})+\mathcal{L}_{aux}(x^{s})$ $\mathcal{L}_{aux}(x^{t})$
Fully Test-Time Adaptation (TTA, Ours) $x^{t}$ $\mathcal{L}_{unsup}(x^{t})$
3 Method
@@ 3.1 OVSS with VLMs
Following recent approaches for OVSS naclip; wang2025sclip, we employ a transformer-based VLM to extract visual and text features from the image and concepts in natural language. Specifically, we feed the image $\mathbf{X}$ into the ViT-based vision encoder to extract a visual token matrix $\mathbf{F}=\big[\mathbf{f}_{[\texttt{cls}]},\mathbf{f}_{1},\dots,\mathbf{f}_{N}\big]$ with each $\mathbf{f}_{i}\in^{D}$, where $N\!=\!\lfloor H/s\rfloor\!\times\!\lfloor W/s\rfloor$ is the number of patches of size $s\!\times\!s$ in the image and [cls] is the CLS token for classification. We define $\mathbf{Q}=\big[\mathbf{q}_{[\texttt{cls}]},\mathbf{q}_{1},\dots,\mathbf{q}_{N}\big]$, with each $\mathbf{q}_{i}\in^{D^{\prime}}$, the output features before the projection layer: $\mathbf{F}=\text{proj}(\mathbf{Q})$. At the same time, the text encoder is employed to extract text features $\mathbf{t}_{k}\in^{D}$ for each concept $C_{k}\in\mathcal{C}$. This is achieved by combining $C_{k}$ with a text prompt template, for instance “A photo of a $C_{k}$]” or “An image of a $C_{k}$]” where $C_{k}$ is an arbitrary text description like “white horse”.
@@
$\arg\max_{k}\ \mathit{sim}\big(\mathbf{f}_{[\texttt{cls}]},\mathbf{t}_{k}\big),\ \ \text{where }\mathit{sim}(\mathbf{x},\mathbf{y})=\frac{\mathbf{x}\cdot\mathbf{y}}{\|\mathbf{x}\|\|\mathbf{y}\|}.$ (1)
@@ 3.2 MLMP: Proposed Method
$p_{ik}\,=\,\frac{\exp\big(\mathit{sim}(\mathbf{f}_{i},\mathbf{t}_{k})/\tau\big)}{\sum_{k^{\prime}=1}^{|\mathcal{C}|}\exp\big(\mathit{sim}(\mathbf{f}_{i},\mathbf{t}_{k^{\prime}})\big/\tau)}.$ (2)
@@
Uncertainty-Aware Multi-Level Fusion.  In VLM-based classification, the CLS token in the last layer of the visual encoder is typically used to compute the class label probabilities. This approach relies on the idea that the relevant information for classification lies at the end of the ViT and that intermediate layers serve to transform features. In segmentation, however, features from intermediate layers are often used to capture complementary information at different scales unet. This hypothesis is validated in Table 2 as well as Figure 3, showing that a higher segmentation mIoU is obtained when combining the features from different intermediate layers.
@@
Image-Level Entropy Minimization.  Since the CLS token is not directly linked to individual patch predictions but rather captures a more global representation of the input, we also include an image-level entropy (ILE) minimization term specifically for this token. As illustrated in Figure 1, the CLS token demonstrates increased robustness and reliability. This term, which encourages the model to produce more confident global predictions is expressed as:
@@
Multi-Prompt Adaptation.  Prior work on TTA for classification watt has shown the usefulness of leveraging multiple prompt templates in VLM to encode class labels, based on the idea that the templates capture complementary information about these classes. As shown in Figure 1a, the sensitivity to the choice of prompt templates is even more pronounced in segmentation tasks, where fine-grained spatial predictions are required. Using multiple templates acts as cross-modal regularization, encouraging more stable and generalized learning signals. While different from image augmentation, it can be seen as a strong, safe, and lightweight text-space augmentation.
Rather than averaging the weights adapted from different prompt templates as in watt—a computationally expensive approach for dense prediction tasks such as segmentation—our method minimizes our proposed UAML and ILE losses across these templates. Let $\mathbf{T}_{t}$ be the text features obtained using the $t$-th template. Our final adaptation loss is defined as:
@@
A proof of Prop. 1 is provided in the Appendix. ∎
@@ 4 Experiments
Datasets.  In traditional TTA for segmentation, two datasets are commonly employed to simulate domain shifts—one for model training and another for adaptation during inference (e.g., GTAV and Cityscapes)—as both must share the same semantic label space. In our study, as this is the first exploration of TTA for VLMs in segmentation tasks and given that VLMs are pre-trained, we draw inspiration from ImageNet-C cifar10c&10.1 to introduce 15 synthetic corruptions on segmentation datasets. Our experiments are conducted on Pascal VOC 20 (v20), Pascal VOC 21 (v21) PascalVoc, Pascal Context 59 (P59), Pascal Context 60 (P60) PascalContext, and Cityscapes Cityscapes, incorporating both original version (clean) and the synthetic 15 corruptions (denoted with a “-C” suffix). For COCO-Stuff CocoStuff and COCO-Object CocoObject, we use only the original versions. To further evaluate robustness under real and rendered distributional shifts, we additionally include ACDC acdc—capturing real-world adverse conditions such as fog, night, rain, and snow—and GTA-V gtav, which provides photorealistic, game-rendered urban scenes. This extended setup results in 87 distinct test scenarios encompassing synthetic, real, and rendered shifts, enabling a comprehensive evaluation of MLMP across diverse conditions.
@@
ViT-L/14 Layer Range
$L_{24}$
$L_{23\text{--}24}$
(last two)
$L_{22\text{--}24}$
(last three)
$L_{19\text{--}24}$
(last 25%)
$L_{13\text{--}24}$
(last 50%)
$L_{7\text{--}24}$
(last 75%)
$L_{1\text{--}24}$
(all layers)
@@ 5.1 Ablation studies
[…]
Visualization of Layer-Wise Confidence Weights.  We visualize the mean and standard deviation of the learned layer weights to better understand the behavior of our uncertainty-aware fusion strategy. As shown in Figure 3, deeper layers tend to receive higher confidence, though earlier layers also contribute, especially under corrupted conditions. This variation is most pronounced in the Cityscapes dataset, where the distribution fluctuates more across layers and corruption types. In contrast, the COCO-Stuff dataset shows a flatter distribution, where the final layer is not consistently the most influential. These results underscore the core strength of our fusion mechanism: its ability to adaptively reweight layers based on input conditions, assigning greater importance to those that remain more reliable under distribution shifts and corruption. Please refer to the Appendix for additional results on other datasets.
Effect of Number of Prompt Templates.  To isolate this effect, we evaluate performance using different numbers of prompt templates while disabling both the multi-level fusion and the image-level entropy (ILE) term in Eq.10. As shown in Figure5, increasing the number of templates improves performance up to 7, after which the gains begin to saturate or slightly decline. This trend holds across both clean and corrupted settings, indicating that a moderate number of diverse prompt templates is sufficient for MLMP. We therefore use 7 templates by default in our main experiments. Please refer to the Appendix for template details.
@@
Full Component Analysis.  Table 3 presents an extensive ablation evaluating the contribution of each component in our MLMP strategy. Each proposed element yields consistent gains when added independently, but it is their combination that delivers the highest overall performance across both clean and corrupted settings, highlighting their strong complementary effects within a unified framework. Additionally, we provide further ablations in the Appendix, including alternative OVSS backbones, different VLMs, ViT architecture variants, computational complexity analysis, and MLMP segmentation map visualizations, as well as discussions on the effect of longer prompts, effect of adaptation iterations, and episodic vs. online adaptation.
Table 4: mIoU comparison of MLMP and baselines across several datasets. CLIPArTT could not be run for a few cases owing to GPU memory shortages. Full per-dataset results are in the Appendix.